One angle in a triagle is 10 degrees more than the smallest angle in the triangle, while the other angle is 20 degrees more than the smallest angle. Find the measures of the three angles.
Answer by rfer(16322)

x+x+10+x+20=180
3x=150
x=50
x+10=60
x+20=70